​Anzio, Inc., has two classes of shares. Class B has 10 times the voting rights as Class A. If you own 18%of the Class A shares and 23% of the Class B​ shares, what percentage of the total voting rights do you​ hold?
Answer #1

Let Class has 1 Voting Right and Class B has 10 Voting Rights

Total Voting Rishts = Class A voting Risghts + Class B Voting Risghts

= 100 + 1000

= 1100

Your Share of Voting Rights = 18 + 23(10)

= 18 + 230

= 248

% of Voting risghts = Share of Voting Rights / Total Rights

= 248 / 1100

= 0.2255 i.e 22.55%

Pls comment, if any further assistance is required.